Computer Modern with T1 encoding

THE QUICK BROWN FOX JUMPS OVER THE LAZY DOG the quick brown fox jumps over the lazy dog ABCDEFGHIJKLMNOPQRSTUVWXYZ abcdefghijklmnopqrstuvwxyz Digits: 0123456789 9876543210 Other ASCII: ‘!@#$%&*() -+=[]{}—:;”’¡¿?,./ Other TEX: ı ȷ ff fi fl ffi ffl Ł ł … “ ” Æ Œ æ œ Accents: À Á Â Ã Ä à á â ã ä ā ȧ ă ǎ a̋ a̧ a. ā Ç ç

BrushScriptX-Italic with T1 encoding

THE QUICK BROWN FOX JUMPS OVER THE LAZY DOG the quick brown fox jumps over the lazy dog ABCDEFGHIJKLMNOPQRSTUVWXYZ abcdefghijklmnopqrstuvwxyz Digits: 0123456789 9876543210 Other ASCII: ‘!@#$%&*()_-+=[]{}|:;"’<>?,./\ Other TEX: ı ȷ ff fi fl ffi ffl Ł ł … “ ” Æ Œ æ œ Accents: À Á Â Ã Ä à á â ã ä ā ȧ ă ǎ a̋ a̧ a. ā Ç ç